OpenStudy (anonymous):

Find the slope-intercept equation for the line with the indicated slope and y-intercept. Slope 7; y-intercept (0,0)

OpenStudy (anonymous):

y-yo = m(x-xo) just replace

OpenStudy (anonymous):

I don't get the formula?

OpenStudy (dan815):



OpenStudy (anonymous):

\[y _{1} - y _{1} = m(x _{1} - x _{2_{}})\] \[y-0 = 7(x-0)\] \[y = 7x\]

OpenStudy (jhannybean):

point slope form follows the equation \[\large y-y_{1}=m(x-x_{1})\] where \(\large (x_{1},y_{1}) = (0,0)\)and \(\large m=7\) Plugging it into our equation, we will have \[\large y-0 = 7(x-0)\] Now we simplify\[\large y= 7x \]